# Pólya’s conjecture for thin products

### v2

@@ Abstract.
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded Euclidean domain. According to the famous Weyl law, both its Dirichlet eigenvalue $\lambda_{k}(\Omega)$ and its Neumann eigenvalue $\mu_{k}(\Omega)$ have the same leading asymptotics $w_{k}(\Omega)=C(d,\Omega)k^{2/d}$ as $k\to\infty$. G. Pólya conjectured in 1954 that each Dirichlet eigenvalue $\lambda_{k}(\Omega)$ is greater than $w_{k}(\Omega)$, while each Neumann eigenvalue $\mu_{k}(\Omega)$ is no more than $w_{k}(\Omega)$. In this paper we prove Pólya’s conjecture for thin products, i.e. domains of the form $(a\Omega_{1})\times\Omega_{2}$, where $\Omega_{1},\Omega_{2}$ are Euclidean domains, and $a$ is small enough. We also prove that the same inequalities hold if $\Omega_{2}$ is replaced by a Riemannian manifold, and thus get Pólya’s conjecture for a class of “thin” Riemannian manifolds with boundary.
Partially supported by National Key R and D Program of China 2020YFA0713100, and by NSFC no. 12171446.
@@ 1. Introduction
by L. Friedlander in [15] (See also N. Filonov [6]), answering a conjecture of L. E. Payne [27]. Starting from H. Weyl ([35]), the asymptotic behavior of the eigenvalues $\lambda_{k}(\Omega)$ and $\mu_{k}(\Omega)$ as $k\to\infty$ has attracted a lot of attention. In fact, both $\lambda_{k}(\Omega)$ and $\mu_{k}(\Omega)$ admit the same leading term asymptotics
@@
where $|\Omega|$ represents the volume of $\Omega$, and $\omega_{d}$ is the volume of the unit ball in $\mathbb{R}^{d}$. In his classical book [29], G. Pólya conjectured (in a slightly weaker form for the Neumann case) that for each $k$, the $k^{\mathrm{th}}$ Dirichlet eigenvalue
@@
As observed by G. Pólya, these conjectured inequalities hold for all rectangles. As for arbitrary domain, the conjecture holds for $k=1$ (the Faber-Krahn inequality ([5], [19]) for the Dirichlet eigenvalue, and the Szegö-Weinberger inequality ([33], [34]) for the Neumann case) and $k=2$ (the Krahn-Szegö inequality ([20]) for the Dirichlet case, and recently proved by D. Bucur and A. Henrot in [4] for the Neumann case). The first major progress on the conjecture was made by G. Pólya himself in 1961 ([30]), in which he presented an elegant proof of his conjecture for planar tiling domains (in fact G. Pólya’s proof for the Neumann eigenvalue case was a bit complicated and assumed the tiling to be regular. The regularity assumption was removed and the proof was simplified in 1966 by R. Kellner [17]). Very recently, N. Filonov, M. Levitin, I. Polterovich and D. Sher ([7]) proved that Pólya’s conjecture holds for planar disks (and for Euclidean balls of all dimensions for the Dirichlet case), and thus gave the first non-tiling planar domain for which Pólya’s conjecture is known to be true. For an arbitrary Euclidean domain $\Omega\subset\mathbb{R}^{d}$, P. Li and S.T. Yau proved in [23] that
(1.4) $\sum_{j=1}^{k}\lambda_{k}(\Omega)\geq\frac{d}{d+2}\frac{4\pi^{2}}{(\omega_{d}|\Omega|)^{\frac{2}{d}}}k^{\frac{d+2}{d}},$
@@
P. Kröger proved in [18] a similar weaker upper bound for the Neumann eigenvalues of any Euclidean domain with piecewise smooth boundary. Another important class of domains satisfying Pólya’s conjecture was obtained by A. Laptev [21], in which he proved that if Pólya’s conjecture (1.2) holds for $\Omega_{1}\subset\mathbb{R}^{d_{1}}$, where $d_{1}\geq 2$, then Pólya’s conjecture (1.2) also holds for any domain of the form $\Omega=\Omega_{1}\times\Omega_{2}$. One key ingredient in his proof is the following inequality (which is a special case of Berezin-Lieb inequality ([3], [24]) and is equivalent to Li-Yau’s inequality (1.4) above) for the Riesz mean,
@@
using which one can get Pólya’s conjecture (1.3) for $\Omega=\Omega_{1}\times\Omega_{2}$ provided $\Omega_{1}$ satisfies (1.3) and has dimension $d_{1}\geq 2$. For other recent progresses concerning Pólya’s conjecture, we refer to [8],[11], [12], [13], [25] etc. In this paper we will prove Pólya’s conjecture for domains of product type that are “thin” in one component, namely regions of the form
@@
for $a$ small enough, without assuming that $\Omega_{1}$ or $\Omega_{2}$ satisfies Pólya’s conjecture. We first prove
@@
Let $\Omega_{1}\subset\mathbb{R}^{d_{1}}$ and $\Omega_{2}\subset\mathbb{R}^{d_{2}}$ be bounded Euclidean domains, where $d_{1},d_{2}\geq 2$, and $\Omega_{2}$ has piecewise smooth boundary.
  1. (1)
@@
    If $\Omega_{1}$ has Lipschitz boundary, then there exists $a_{0}>0$ (depends on $\Omega_{1}$ and $\Omega_{2}$) such that for any $0<a<a_{0}$, the product $\Omega=a\Omega_{1}\times\Omega_{2}$ satisfies the Dirichlet Pólya’s conjecture (1.2).
  2. (2)
    If $\Omega_{1}$ has $C^{1}$ boundary, then there exists $a_{0}>0$ (depends on $\Omega_{1}$ and $\Omega_{2}$) such that for any $0<a<a_{0}$, the product $\Omega=a\Omega_{1}\times\Omega_{2}$ satisfies the Neumann Pólya’s conjecture (1.3).
@@
Here is the strategy of proof: Following Laptev’s argument [21], we write the eigenvalue counting function of $a\Omega_{1}\times\Omega_{2}$ as the sum of many eigenvalue counting functions of $\Omega_{2}$. Although we don’t have Pólya’s inequality for $\Omega_{2}$, we do have weaker inequalities (See (2.6) and (2.7) below) that follow from Seeley’s version of the two-term Weyl law (which only requires $\Omega_{2}$ to have piecewise smooth boundary). Now instead of applying Laptev’s inequalities on Riesz mean above, we apply stronger two-term inequalities on Riesz mean, namely (2.8) obtained by R. Frank and S. Larson in [10] (for $\Omega_{1}$ to have Lipschitz boundary) and its Neumann analogue, (2.9), obtained by R. Frank and L. Geisinger in [9] (for $\Omega_{1}$ to have $C^{1}$ boundary) to control the sum of both terms in Seeley’s inequalities. By comparing what we lose from Seeley’s two-term bound and what we gain from these two-term Riesz mean bound, we are able to prove that for $a$ small enough, Pólya’s inequalities hold for $\lambda$ large enough (which depends on $a$). For smaller $\lambda$, we use a simple observation that Laptev’s inequalities above are in fact strict (See §2.2 below), and thus (by taking $a$ even smaller) give us the demanded gap to prove Pólya’s inequality. This argument works perfectly well for $d_{2}\geq 3$, but the first part fails for $d_{2}=2$ since in this case we can’t apply Laptev-type inequality on Riesz mean (which requires $\gamma=\frac{d_{2}-1}{2}\geq 1$) to control the sum of the second term of Seeley’s inequality. Fortunately, we can overcome this problem by using Li-Yau’s estimate (1.5) above and an explicit integral computation. The Neumann case is a bit simpler, since we only need Weyl’s law to control the second term in Seeley’s inequality for large $\lambda$, and thus we don’t need to distinguish the case $d_{2}=2$ with $d_{2}\geq 3$.
@@
Note that Laptev’s argument does not work for the case $d_{1}=1$, since the inequalities (1.6) and (1.8) require $\gamma\geq 1$. Even though the interval $(0,1)$ tiles $\mathbb{R}$, it is still not known whether $(0,1)\times\Omega$ satisfies Pólya’s conjecture for general $\Omega$. In the second part of this paper, we turn to study Pólya’s conjecture for thin products $(0,a)\times\Omega$. Instead of writing the eigenvalue counting function of $(0,a)\times\Omega$ as the sum of many eigenvalue counting functions of $(0,a)$ (which is a tiling domain) that we have a nice control, we will write it as the sum of many eigenvalue counting functions of $\Omega$ and apply Seeley’s two-term inequalities. By carefully analyzing the two sums, we shall prove that in this case, all thin products satisfy Pólya’s conjecture:
@@
Since scaling will not affect Pólya’s inequalities, we immediately see that for any bounded Euclidean domain $\Omega$, there exists a constant $C>0$ such that for all $A>C$, $(0,1)\times A\Omega$ satisfies Pólya’s conjecture. Unfortunately we still can’t prove Pólya’s conjecture for products of the form $(0,1)\times a\Omega$ for small $a$, which obviously implies Pólya’s conjecture for $(0,1)\times\Omega$. The last part of this paper devoted to Pólya’s inequalities for Riemannian manifolds with boundary. Although the original conjecture was only proposed for Euclidean domains, people do study the analogous problem in the more general Riemannian setting. For example, P. Bérard and G. Besson proved in [2] that for a 2-dimensional hemisphere (or a quarter of a sphere, or even an octant of a sphere), both Dirichlet eigenvalues and Neumann eigenvalues satisfy Pólya’s inequalities above. Recently in [14], P. Freitas, J. Mao and I. Salavessa studied the problem for hemispheres in arbitrary dimension. They showed that (1.3) holds for Neumann eigenvalues of hemispheres in any dimension, while (1.2) fails for Dirichlet eigenvalues when $d>2$, and they derived sharp inequality for Dirichlet eigenvalues by adding a correction term. It is thus a natural problem to find out more Riemannian manifolds with boundary satisfying Pólya’s inequalities. Note that in the proof of Theorem 1.1 and Theorem 1.2, for $\Omega_{2}$ and $\Omega$ we mainly used Seeley’s two-term Weyl’s inequality. As a result, by literally repeating the proof one can easily see that for any closed Riemannian manifold $M$, the Neumann eigenvalues of the product $a\Omega\times M$ satisfy Pólya conjecture (1.3) as long as $a$ is small enough. For the Dirichlet case, there will be one extra term (since 00 is an eigenvalue of $M$) in the eigenvalue counting function of the product, namely the number of eigenvalues of $\Omega$ that is less than $a^{2}\lambda$, which can be explicitly calculated if $d_{1}=\dim\Omega=1$ and can be controlled via Li-Yau’s estimate (1.5) if $d_{2}\geq 2$. As a result, we are able to prove that Pólya’s conjecture holds for such Riemannian manifolds with boundary:
@@
Let $\Omega\subset\mathbb{R}^{d_{1}}$ be a bounded domain with $C^{1}$ boundary and $(M,g)$ be a closed Riemannian manifold of dimension $d_{2}\geq 2$. Then there exists $a_{0}>0$ (depends on $\Omega$ and $M$) such that for any $0<a<a_{0}$, $a\Omega\times M$ satisfies Pólya’s conjecture (1.2) and (1.3).
The arrangement of this paper is as follows. In Section 2 we will list the two-term inequalities for the eigenvalues counting functions and for the Riesz means that will be used later, and also prove strict version of Laptev’s inequalities for completeness. In Section 3 we will prove Theorem 1.1, and in Section 4 we will prove Theorem 1.2. In Section 5 we will turn to the Riemannian manifold setting and prove Theorem 1.3. Moreover we will explain how to get similar results for a larger class of eigenvalue problems. Finally in Section 6 we will give an explicit non-tiling planar domain $\Omega$ and explicitly calculate the constant involved in the proof, and as a result, show that the Dirichlet eigenvalues of $[0,\frac{1}{4\pi}]\times\Omega$ for that $\Omega$ satisfies (1.2). Similarly in the Riemannian setting we show that $(0,a)\times S^{2}$ satisfies Pólya’s conjecture for $a\leq\frac{\pi}{24}$, but will break both Pólya’s inequalities if $a$ is large.

### v3

@@ Abstract.
Let $\Omega\subset\mathbb{R}^{d}$ be a bounded Euclidean domain. According to the famous Weyl law, both its Dirichlet eigenvalue $\lambda_{k}(\Omega)$ and its Neumann eigenvalue $\mu_{k}(\Omega)$ have the same leading asymptotics $w_{k}(\Omega)=C(d,|\Omega|)k^{2/d}$ as $k\to\infty$. G. Pólya conjectured in 1954 that each Dirichlet eigenvalue $\lambda_{k}(\Omega)$ is greater than $w_{k}(\Omega)$, while each Neumann eigenvalue $\mu_{k}(\Omega)$ is no more than $w_{k}(\Omega)$. In this paper we prove Pólya’s conjecture for thin products, i.e. domains of the form $(a\Omega_{1})\times\Omega_{2}$, where $\Omega_{1},\Omega_{2}$ are Euclidean domains, and $a$ is small enough. We also prove that the same inequalities hold if $\Omega_{2}$ is replaced by a Riemannian manifold, and thus get Pólya’s conjecture for a class of “thin” Riemannian manifolds with boundary.
@@ 1. Introduction
by L. Friedlander in [18] (See also N. Filonov [6]), answering a conjecture of L. E. Payne [36].
Starting from H. Weyl ([46]), the asymptotic behavior of the eigenvalues $\lambda_{k}(\Omega)$ and $\mu_{k}(\Omega)$ as $k\to\infty$ has attracted a lot of attention. In fact, both $\lambda_{k}(\Omega)$ and $\mu_{k}(\Omega)$ admit the same leading term asymptotics
@@
where $|\Omega|$ represents the volume of $\Omega$, and $\omega_{d}$ is the volume of the unit ball in $\mathbb{R}^{d}$.
In his classical book [38], G. Pólya conjectured (in a slightly weaker form for the Neumann case) that for each $k$, the $k^{\mathrm{th}}$ Dirichlet eigenvalue
@@
As observed by G. Pólya, these conjectured inequalities hold for all rectangles. For arbitrary domain, the conjecture holds for $k=1$ (the Faber-Krahn inequality ([5], [24]) for the Dirichlet eigenvalue, and the Szegö-Weinberger inequality ([43], [45]) for the Neumann case) and $k=2$ (the Krahn-Szegö inequality ([25]) for the Dirichlet case, and recently proved by D. Bucur and A. Henrot in [4] for the Neumann case).
The first major progress on the conjecture was made by G. Pólya himself in 1961 ([39]), in which he presented an elegant proof of his conjecture for planar tiling domains (in fact G. Pólya’s proof for the Neumann eigenvalue case relied on the assumption of regular tiling, which was removed by R. Kellner in 1966 [22]). The idea is to compare the $k$th eigenvalue of $\Omega$ to the $kn_{r}$th eigenvalue of the unit square, where $n_{r}$ is the number of $r\Omega$’s that almost tile the unit square, and then apply Weyl’s asymptotics to the later.
For an arbitrary Euclidean domain $\Omega\subset\mathbb{R}^{d}$, P. Li and S.T. Yau proved in [30] that
(1.4) $\sum_{j=1}^{k}\lambda_{j}(\Omega)\geq\frac{d}{d+2}\frac{4\pi^{2}}{(\omega_{d}|\Omega|)^{\frac{2}{d}}}k^{\frac{d+2}{d}},$
@@
In [23], P. Kröger established two upper bounds for the Neumann eigenvalues of any Euclidean domain with piecewise smooth boundary:
$\sum_{j=1}^{k-1}\mu_{j}(\Omega)\leq\frac{d}{d+2}\frac{4\pi^{2}}{(\omega_{d}|\Omega|)^{\frac{2}{d}}}k^{\frac{d+2}{d}},$
and
(1.6) $\mu_{k}(\Omega)\leq\bigg{(}\frac{d+2}{2}\bigg{)}^{\frac{2}{d}}\frac{4\pi^{2}}{(\omega_{d}|\Omega|)^{\frac{2}{d}}}k^{\frac{2}{d}}.$
Very recently, N. Filonov improved the bound (1.6) for convex bounded domains in $\mathbb{R}^{2}$ (see [9]), obtaining a result that is closer to the upper bound predicted by Pólya’s conjecture.
Another important class of domains satisfying Pólya’s conjecture was obtained by A. Laptev [26], in which he proved that if Pólya’s conjecture (1.2) holds for $\Omega_{1}\subset\mathbb{R}^{d_{1}}$, where $d_{1}\geq 2$, then Pólya’s conjecture (1.2) also holds for any domain of the form $\Omega=\Omega_{1}\times\Omega_{2}$. One key ingredient in his proof is the following inequality (which is a special case of Berezin-Lieb inequality ([3], [32]) and is equivalent to Li-Yau’s inequality (1.4) above) for the Riesz mean,
@@
using which one can get Pólya’s conjecture (1.3) for $\Omega=\Omega_{1}\times\Omega_{2}$ provided $\Omega_{1}$ satisfies (1.3) and has dimension $d_{1}\geq 2$. For other recent progresses concerning Pólya’s conjecture, we refer to [14], [15], [16], [33] etc.
Recently, by developing the links between Laplacian eigenvalues of planar disks with certain lattice counting problems, N. Filonov, M. Levitin, I. Polterovich and D. Sher ([7]) proved that Pólya’s conjecture holds for planar disks (and for Euclidean balls of all dimensions for the Dirichlet case), and thus gave the first non-tiling planar domain for which Pólya’s conjecture is known to be true. A key ingredient in the proof is certain uniform bounds between the eigenvalue and lattice point counting functions. For the Neumann case, they apply different tricks to handle large eigenvalues and small eigenvalues. Building upon and extending the methods developed for disks and balls, very recently, they established the validity of Pólya’s conjecture (1.2) for annular domains (see [8]).
In this paper we will prove Pólya’s conjecture for domains of product type that are “thin” in one component, namely regions of the form
@@
for $a$ small enough, without assuming that $\Omega_{1}$ or $\Omega_{2}$ satisfies Pólya’s conjecture. In particular, we obtain lots of non-tiling domains satisfying Pólya’s conjecture. In the proofs we combine tricks used in [7], [26] and [39]. More precisely, we treat large eigenvalues and small eigenvalues separately, we use Weyl law extensively for large eigenvalues, and the product structure lies in the core of the proof.
We first prove
@@
Let $\Omega_{1}\subset\mathbb{R}^{d_{1}}$ and $\Omega_{2}\subset\mathbb{R}^{d_{2}}$ be bounded Euclidean domains, where $d_{1},d_{2}\geq 2$, $\Omega_{1}$ has Lipschitz boundary and $\Omega_{2}$ has piecewise smooth boundary. Then, there exists $a_{0}>0$ (depends on $\Omega_{1}$ and $\Omega_{2}$) such that for any $0<a<a_{0}$, the product $\Omega=a\Omega_{1}\times\Omega_{2}$ satisfies Pólya’s conjecture (1.2) and (1.3).
The dependence of the constant $a_{0}$ with respect to $\Omega_{1}$ and $\Omega_{2}$ arises from various constants in the proof (including the constants needed for the two-term Riesz mean inequalities and Seeley’s inequalities below), and thus are subtle in general. However, if $\Omega_{1}$ is convex, then the dependence on $\Omega_{1}$ is quite explicit (in terms of its diameter, in-radius, volume and surface area). See Remark 3.1 and Remark 3.2 below. We will give a class of domains for which the dependence on $\Omega_{2}$ is explicitly computable at the end of this paper, see Remark 7.2.
@@
Here is the strategy of proof: Following Laptev’s argument [26], we write the eigenvalue counting function of $a\Omega_{1}\times\Omega_{2}$ as the sum of many eigenvalue counting functions of $\Omega_{2}$. Although we don’t have Pólya’s inequality for $\Omega_{2}$, we do have weaker inequalities (See (2.7) and (2.8) below) that follow from Seeley’s version of the two-term Weyl law (which only requires $\Omega_{2}$ to have piecewise smooth boundary). Now instead of applying Laptev’s Berezin inequalities on Riesz mean above, we apply stronger two-term inequalities on Riesz mean, namely (2.9) and (2.10) obtained by R. Frank and S. Larson in [12] (see also [10], [11]) to control the sum of the first term in Seeley’s inequalities. We will have to distinguish the two boundary conditions:
  • In the Dirichlet setting, we also use Laptev’s Riesz sum inequality to control the sum of the second term in Seeley’s inequalities. By comparing what we lose from Seeley’s two-term bound and what we gain from these two-term Riesz mean bound, we are able to prove that for $a$ small enough, Pólya’s inequalities hold for $\lambda$ large enough (which depends on $a$). For smaller $\lambda$, we use Proposition 2.1 in [13], and thus (by taking $a$ even smaller) give us the demanded gap to prove Pólya’s inequality. This argument works perfectly well for $d_{2}\geq 3$, but fails for $d_{2}=2$ since we can’t apply Laptev-type inequality on Riesz mean (which requires $\gamma=\frac{d_{2}-1}{2}\geq 1$) to control the sum of the second term of Seeley’s inequality. Fortunately, we can overcome this problem by using Li-Yau’s estimate (1.5) above and an explicit integral computation.
  • In the Neumann setting, one can’t use the same argument since we also need an upper bound of (the sum of) the second term in Seeley’s inequality for large $\lambda$, which does not follow from any Riesz mean inequality for Neumann eigenvalues. So instead we use Weyl’s law directly to control the second term, and as a result we don’t need to distinguish the case $d_{2}=2$ with $d_{2}\geq 3$. Another difference with the Dirichlet case is that we do have very small eigenvalues in this case, but fortunately the classical Szegö-Weinberger inequality is enough for us to handle these eigenvalues.
@@
Note that Laptev’s argument does not work for the case $d_{1}=1$, since the inequalities (1.7) and (1.9) require $\gamma\geq 1$. Even though the interval $(0,1)$ tiles $\mathbb{R}$, it is still not known whether $(0,1)\times\Omega$ satisfies Pólya’s conjecture for general $\Omega$. In the second part of this paper, we turn to study Pólya’s conjecture for thin products $(0,a)\times\Omega$. Instead of writing the eigenvalue counting function of $(0,a)\times\Omega$ as the sum of many eigenvalue counting functions of $(0,a)$ (which is a tiling domain) that we have a nice control, we will write it as the sum of many eigenvalue counting functions of $\Omega$ and apply Seeley’s two-term inequalities. We then carefully analyze the two sums and show that the sums are controlled by some explicit integrals (similar trick was used in [7]). As a result we shall prove that in this case, all thin products satisfy Pólya’s conjecture:
@@
Since scaling will not affect Pólya’s inequalities, we immediately see that for any bounded Euclidean domain $\Omega$, there exists a constant $C>0$ such that for all $A>C$, $(0,1)\times A\Omega$ satisfies Pólya’s conjecture. Unfortunately we still can’t prove Pólya’s conjecture for products of the form $(0,1)\times a\Omega$ for small $a$, which obviously implies Pólya’s conjecture for $(0,1)\times\Omega$.
The next part of this paper devotes to Pólya’s inequalities for Riemannian manifolds with boundary. Although the original conjecture was proposed only for Euclidean domains, people did study the analogous problem in the more general Riemannian setting. For example, P. Bérard and G. Besson proved in [2] that for a 2-dimensional hemisphere (or a quarter of a sphere, or even an octant of a sphere), both Dirichlet eigenvalues and Neumann eigenvalues satisfy Pólya’s inequalities above. Recently in [17], P. Freitas, J. Mao and I. Salavessa studied the problem for hemispheres in arbitrary dimension. They showed that (1.3) holds for Neumann eigenvalues of hemispheres in any dimension, while (1.2) fails for Dirichlet eigenvalues when $d>2$, and they derived sharp inequality for Dirichlet eigenvalues by adding a correction term.
It is thus a natural problem to find out more Riemannian manifolds with boundary satisfying Pólya’s inequalities. Note that in the proof of Theorem 1.1 and Theorem 1.2, for $\Omega_{2}$ and $\Omega$ we mainly used Seeley’s two-term Weyl’s inequality. As a result, by literally repeating the proof one can easily see that for any closed Riemannian manifold $M$, the Neumann eigenvalues of the product $a\Omega\times M$ satisfy Pólya conjecture (1.3) as long as $a$ is small enough. For the Dirichlet case, there will be one extra term (since 00 is an eigenvalue of $M$) in the eigenvalue counting function of the product, namely the number of eigenvalues of $\Omega$ that is less than $a^{2}\lambda$, which can be explicitly calculated if $d_{1}=\dim\Omega=1$ and can be controlled via Li-Yau’s estimate (1.5) if $d_{2}\geq 2$. As a result, we are able to prove that Pólya’s conjecture holds for such Riemannian manifolds with boundary:
@@
Let $\Omega\subset\mathbb{R}^{d_{1}}$ be a bounded domain with Lipschitz boundary and $(M,g)$ be a closed Riemannian manifold of dimension $d_{2}\geq 2$. Then there exists $a_{0}>0$ (depends on $\Omega$ and $M$) such that for any $0<a<a_{0}$, $a\Omega\times M$ satisfies Pólya’s inequalities (1.2) and (1.3).
Another natural question is: if $\Omega$ satisfies Pólya’s conjecture, and $\Omega^{\prime}$ is “sufficiently close” to $\Omega$ in some sense, can we prove Pólya’s conjecture for $\Omega^{\prime}$? Applying the techniques we developed in the proofs of Theorem 1.1 and 1.2, we will give a positive result in a special product setting. More precisely, we will show that if $\Omega_{2}\subset\mathbb{R}^{d_{2}}$ ($d_{2}\geq 2$) satisfies Pólya’s conjecture, then for any $\Omega_{3}\subset\Omega_{2}$, the product domain $\Omega_{1}\times(\Omega_{2}\setminus a\Omega_{3})$ (which is not a thin product, but the complement of a thin product) satisfies Pólya’s conjecture for $a$ small enough. See Theorem 6.1 and Theorem 6.2 for precise statement.
The arrangement of this paper is as follows. In Section 2 we will list the two-term inequalities for the eigenvalues counting functions and for the Riesz means that will be used later. In Section 3 we will prove Theorem 1.1, and in Section 4 we will prove Theorem 1.2. In Section 5 we will turn to the Riemannian manifold setting and prove Theorem 1.3. Moreover we will explain how to get similar results for a larger class of eigenvalue problems. In Section 6, we prove Pólya’s conjecture for $\Omega_{1}\times(\Omega_{2}\setminus a\Omega_{3})$, where $\Omega_{1}\times\Omega_{2}$ is the product domain in Laptev’s theorem. Finally in Section 7 we will give an explicit non-tiling planar domain $\Omega$ and explicitly calculate the constant involved in the proof, and as a result, show that the Dirichlet eigenvalues of $[0,\frac{1}{4\pi}]\times\Omega$ for that $\Omega$ satisfies (1.2).